OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. Christy’s plan for a rectangular dog pen is too small. The perimeter needs to be more than 118 feet. If the width should be nine feet shorter than the length, what should the length be?

OpenStudy (anonymous):

P = 2L + 2w w = L -9 118 = 2L +2(L - 9) 118 = 2L + 2L - 18 136 = 4L L = 34